# Transformers Provably Solve Parity Efficiently with Chain of Thought

### v1

@@ 3.3 CoT without Teacher Forcing
This gives the same order of separation from Theorem 2(2) as in Section 3.2. Hence transformers can learn parities even without teacher forcing, if the consistency of the chain of reasoning is suitably controlled for. Moreover, our result shows that logarithmic time suffices to learn parity by exploiting the hierarchical decomposition in Figure 1. This extends the circuit complexity result in Merrill and Sabharwal (2024), which states that bounded-depth transformers with a logarithmic number of CoT steps can express problems in log-space; Theorem 7 guarantees that transformers of depth one can learn by gradient descent any such function in the exponentially large class $P$.
@@
Sketch of proof. The idea is to inductively show that each $2$-parity subtask $x_{m}$ at level $\ell$ will become solved at time $t=\ell$. When $t\leq\ell-2$, $x_{m}$ will not have access to its child nodes $x_{\mathsf{c}_{1}[m]},x_{\mathsf{c}_{2}[m]}$ since they will also not be optimized, so the weights do not change. At time $\ell-1$, its child nodes learn to output their respective parities with high precision, so the objective is approximately equivalent to that of Theorem 5. Then the gradient signal will similarly concentrate on $w_{\mathsf{c}_{1}[m],m},w_{\mathsf{c}_{2}[m],m}$ and $x_{m}$ will become solved in the next step. It remains to bound the gradients arising from the loss terms further down the chain $\bm{x}_{d+1}\to\cdots\to\bm{x}_{d+k-1}$ (propagation of error), and verify that irrelevant weights $w_{j,m}$ ($\mathsf{p}[j]\neq m$) and already optimized weights do not change. This is done by exploiting the effect of the filter $\iota$ and concentration bounds for the original and augmented samples. ∎
@@ 4 Conclusion
In this paper, we provide a theoretical foundation for training transformers with CoT to perform stepwise reasoning through task decomposition by studying the $k$-parity problem. Our results show that gradient-based learning of parity requires significant iterations without intermediate supervision, but task decomposition using teacher forcing enables efficient learning in a single gradient update. Furthermore, when transformers are trained to generate reasoning chains end-to-end, data augmentation and self-consistency checks enhance their ability to solve complex tasks. Our work offers a foundational understanding of how CoT can be leveraged in fine-tuning transformers to improve multi-step reasoning capability.
@@ Acknowledgments
JK was partially supported by JST CREST (JPMJCR2015). TS was partially supported by JSPS KAKENHI (24K02905) and JST CREST (JPMJCR2115).
@@ Appendix A Proof of Theorem 2
where $Q\subseteq P$ denotes the corresponding subset of the hypothesis space. Note that the argument can be extended to any randomized algorithm and random initialization in a straightforward manner by lifting to the product probability space, and so we consider $Q$ to be fixed. Then for any target parity $p\in Q$, the output $\mathcal{A}(\theta)$ of the algorithm after $T$ steps does not depend on $p$, so the predictor $f=f_{\mathcal{A}(\theta)}$ is also fixed.
@@
$\mathbb{E}_{p\in P}\left[\sup_{\bm{x}}\left\lvert p(\bm{x})-f_{\mathcal{A}(\theta)}(\bm{x})\right\rvert\right]$ $\geq\mathbb{E}_{p\in P}\left[1_{\{p\in Q\}}\sup_{\bm{x}}\left\lvert p(\bm{x})-f(\bm{x})\right\rvert\right]$
@@
$\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\mathcal{A}(\theta)}(\bm{x}))^{2}\right]\geq\mathbb{E}_{p\in P,\bm{x}}\left[1_{\{p\in Q\}}(p(\bm{x})-f(\bm{x}))^{2}\right].$
@@
$\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\mathcal{A}(\theta)}(\bm{x}))^{2}\right]$ $\geq\mathbb{E}_{p\in P,\bm{x}}\left[1_{\{p\in Q\}}(p(\bm{x})-\bar{f}(\bm{x}))^{2}\right]$
@@ Appendix C Proof of Theorem 7
$w_{j,m}^{(t)}-\eta\widetilde{\nabla}_{w_{j,m}}L(\mathbf{W}^{(t)})=2c\eta_{0}\frac{d^{2+\epsilon/16}}{d_{\mathsf{h}[m]-1}^{2}}1_{\{\mathsf{p}[j]=m\}}+O(d^{-\epsilon/16}).$
@@
~$\widetilde{\nabla}_{w_{j,m}}L(\mathbf{W}^{(t)})=O(d\exp(-Cd^{\epsilon/16}))+O(d^{-2-\epsilon/8}),$

### v2

@@ 3.3 CoT without Teacher Forcing
This gives the same order of separation from Theorem 2(2) as in Section 3.2. Hence transformers can learn parities even without teacher forcing, if the consistency of the chain of reasoning is suitably controlled for. Moreover, our result shows that logarithmic time suffices to learn parity by exploiting the hierarchical decomposition in Figure 1. This extends the circuit complexity result in Merrill & Sabharwal (2024), which states that bounded-depth transformers with a logarithmic number of CoT steps can express problems in log-space; Theorem 7 guarantees that transformers of depth one can learn by gradient descent any such function in the exponentially large class $P$.
@@
Sketch of proof. The idea is to inductively show that each $2$-parity subtask $x_{m}$ at level $\ell$ will become solved at time $t=\ell$. When $t\leq\ell-2$, $x_{m}$ cannot utilize its child nodes $x_{\mathsf{c}_{1}[m]},x_{\mathsf{c}_{2}[m]}$ since they will also not be optimized, so the weights do not change. At time $\ell-1$, its child nodes learn to output their parities with high precision, so the objective is approximately equivalent to that of Theorem 5. Then the gradient signal will similarly concentrate on $w_{\mathsf{c}_{1}[m],m},w_{\mathsf{c}_{2}[m],m}$ and $x_{m}$ will become solved in the next step. It remains to bound the gradients arising from the loss terms further down the chain $\bm{x}_{d+1}\to\cdots\to\bm{x}_{d+k-1}$ (propagation of error), and verify that irrelevant weights $w_{j,m}$ ($\mathsf{p}[j]\neq m$) and already optimized weights do not change. The full proof is provided in Appendix C. ∎
4 Numerical Experiments
In this section, we present numerical experiments which support and complement our theoretical findings. Compared to the carefully calibrated step sizes and weight updates in Theorems 5 and 7, these experiments study a more realistic training scenario by taking relatively small learning rates and tracking the loss trajectories over a longer period of training. We train one-layer transformers based on the architecture described in Section 2 to solve a random $k$-parity problem with $64$-bit inputs for $k=8,16,32$. Specifically, we implement and compare the following four models.
  • Direct: $\operatorname{TF}(\cdot)$ is applied to itself $k-1$ times to generate the reasoning chain end-to-end and the model prediction $\hat{\bm{y}}$ is directly compared to the ground truth $\bm{y}$ with the prediction loss $\frac{1}{2n}\lVert\hat{\bm{y}}-\bm{y}\rVert^{2}$.
  • CoT: $\operatorname{TF}(\cdot)$ is applied to itself to generate the reasoning chain end-to-end and the sequence of intermediate states is compared to the ground truth as in (2). Here, we also implement the causal mask in Figure 3 (right) so that only $\log_{2}k$ iterations are needed, for additional stability.
  • CoT + teacher forcing: implements the model in Section 3.2 with teacher forcing.
  • CoT + self-consistency: implements the model in Section 3.3 with the causal mask in Figure 3 (right) and data augmentation for consistency checks. Weight quantization is omitted.
All models are optimized using full-batch gradient descent on 100K $64$-bit samples with a single Tesla T4 GPU. The three CoT models are trained with the ‘CoT loss’ (2) scaled by $\frac{1}{k-1}$ to match the prediction loss of the direct model. Figure 4 shows training curves for the CoT loss (left) and the prediction loss (right) over 350 epochs when $k=32$; results for all $k$ and more details are provided in Appendix D.
We first note that the direct model (red) completely fails to learn the target, plateauing almost immediately. We observed that the weights become nearly uniform so that $\hat{\bm{y}}\approx\bm{0}_{n}$ and the prediction error is stuck at 0.5. This was not improved by using a multilayer transformer instead of repeated composition. The basic CoT model (yellow) is able to significantly decrease CoT loss but fails to fully solve the problem and eventually becomes unstable. Moreover, the prediction loss never improves beyond 0.5. Indeed, due to the hierarchical structure of parity, the model has no chance of making an informative prediction at the last level $\bm{x}_{d+k-1}$ unless all preceding levels have been fully solved. In contrast, we verify that CoT with teacher forcing (blue) solves parity efficiently as predicted in Section 3.2, even with a small learning rate. After a burn-in phase, the CoT loss steadily decreases to nearly zero, at which point the prediction loss also decreases rapidly as the final level is solved.
CoT with self-consistency (green) is also able to solve parity efficiently as predicted. Furthermore, the corresponding CoT loss curve clearly exhibits multiple learning stages. In the beginning, the model is essentially optimizing only the first level as subsequent outputs are zeroed out. After a short burn-in phase, the weights are optimized so that the softmax scores concentrate on the relevant nodes, at which point the CoT loss sharply decreases and the filters for the next level are deactivated, unlocking the next learning stage. This phased optimization repeats until all levels are fully solved and is crucial to arriving at the correct answer (in essence, teacher forcing is doing this at all levels simultaneously). Notably, a similar behavior seems to arise in the basic CoT model as well but fails due to accumulating error, further justifying the use of the filtering mechanism.
These results confirm that training explicitly for CoT generation can improve performance on multi-step tasks, and that controlling error accumulation via teacher forcing or self-consistency is key to ensuring proper step-by-step learning.
Figure 4: CoT loss (left) and prediction loss (right) curves for the four models when $d=64$, $k=32$. For the CoT+consistency model, dashed lines indicate when the filters of each level are deactivated.
@@ 5 Conclusion
In this paper, by focusing on the $k$-parity problem, we provide an initial theoretical foundation for training transformers with CoT to perform stepwise reasoning. Our results show that gradient-based learning of parity requires significant iterations without intermediate supervision, but task decomposition using teacher forcing enables efficient learning in a single gradient update. Furthermore, when transformers are trained to generate reasoning chains end-to-end, data augmentation and self-consistency checks can enhance their ability to solve complex tasks. Our work takes the first steps towards understanding how CoT can be leveraged to improve multi-step reasoning capability of foundation models.
@@ Acknowledgments
JK was partially supported by JST CREST (JPMJCR2015). TS was partially supported by JSPS KAKENHI (24K02905, 20H00576) and JST CREST (JPMJCR2115).
@@ Appendix A Proof of Theorem 2
where $Q\subseteq P$ denotes the corresponding subset of the hypothesis space. Note that the argument can be extended to any randomized algorithm and random initialization in a straightforward manner by lifting to the product probability space, and so we consider $Q$ to be fixed. Then for any target parity $p\in Q$, the output $\theta(\mathcal{A})$ of the algorithm after $T$ steps does not depend on $p$, so the predictor $f=f_{\theta(\mathcal{A})}$ is also fixed.
@@
$\mathbb{E}_{p\in P}\left[\sup_{\bm{x}}\left\lvert p(\bm{x})-f_{\theta(\mathcal{A})}(\bm{x})\right\rvert\right]$ $\geq\mathbb{E}_{p\in P}\left[1_{\{p\in Q\}}\sup_{\bm{x}}\left\lvert p(\bm{x})-f(\bm{x})\right\rvert\right]$
@@
$\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\theta(\mathcal{A})}(\bm{x}))^{2}\right]\geq\mathbb{E}_{p\in P,\bm{x}}\left[1_{\{p\in Q\}}(p(\bm{x})-f(\bm{x}))^{2}\right].$
@@
$\mathbb{E}_{p\in P,\bm{x}}\left[(p(\bm{x})-f_{\theta(\mathcal{A})}(\bm{x}))^{2}\right]$ $\geq\mathbb{E}_{p\in P,\bm{x}}\left[1_{\{p\in Q\}}(p(\bm{x})-\bar{f}(\bm{x}))^{2}\right]$
@@ Appendix C Proof of Theorem 7
$w_{j,m}^{(t)}-\eta\widetilde{\nabla}_{w_{j,m}}L(\mathbf{W}^{(t)},\mathbf{U})=2c\eta_{0}\frac{d^{2+\epsilon/16}}{d_{\mathsf{h}[m]-1}^{2}}1_{\{\mathsf{p}[j]=m\}}+O(d^{-\epsilon/16}).$
@@
~$\widetilde{\nabla}_{w_{j,m}}L(\mathbf{W}^{(t)},\mathbf{U})=O(d\exp(-Cd^{\epsilon/16}))+O(d^{-2-\epsilon/8}),$
@@
Appendix D Experimental Details
For the transformer architecture, the feedforward layer was fixed to the following piecewise quadratic link function:
$\phi(t)=\begin{cases}-4t^{2}-8t-3&t\in[-1,-0.5)\\ 4t^{2}-1&t\in[-0.5,0.5)\\ -4t^{2}+8t-3&t\in[0.5,1].\end{cases}$
For all CoT models, learning rates were fixed to $\eta=15,50,100$ for $k=8,16,32$. For the direct model, the learning rate was scaled to $0.01\eta$ to ensure stability of training. For the self-consistency model, filtering was done through an equivalent weight-based filter, which checks if any softmax score exceeds a threshold value, here set to 0.4. Moreover, we found that adding a 10% fraction of the gradient from the prediction loss to that of the CoT loss resulted in more stable training.
Figure 5 shows CoT loss and prediction loss curves for $k=8,16,32$, extending Figure 4. The direct model fails to learn parity in all cases, while CoT with teacher forcing always learns efficiently. For CoT with self-consistency, a similar analysis as in Section 4 can be applied for $k=8,16$ with two or three distinct learning stages. We also observe that the basic CoT model manages to fully solve the problem for $k=8$ (two intermediate levels) but not for $k=16,32$ (three and four intermediate levels), indicating that assistance (teacher forcing or self-consistency checking) becomes necessary for more complex tasks.
Figure 5: CoT loss (left) and prediction loss (right) curves for the four models when $d=64$, $k=32$ (top), $k=16$ (middle) and $k=8$ (bottom). For the CoT+consistency model, dashed lines indicate when the filters of each level are deactivated.